OpenStudy (anonymous):

A skydiver leaps from a plane, and after a while is accelerating at 7.5 m/s2 downward. If the skydiver has a mass of 61 kg, find the force of air resistance acting on the skydiver.

OpenStudy (michele_laino):

on the skydiver are acting two forces, namely: - the weight force \(W\) - the resistance \(F\) so we can write this formula: \[\Large F = W - ma\] where \(a=7.5\) Using numeric data, we have: \[\Large \begin{gathered} F = W - ma = mg - ma = m\left( {g - a} \right) = \hfill \\ \hfill \\ = 61 \cdot \left( {9.81 - 7.5} \right) = ...newtons \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

of course, the air resistance \(F\) is directed downwards